A father is 10 years older than three times his son's age.
The sum of their ages is 58. What is the difference between the father's age and his son's?

Answer by hellen(1)

Son's age is x.
Dad's age would be 3x+10 because he is 10 years older than 3 times the son's age.
The total age is 58.
Therefore: x+3x+10=58
4x+10=58 (add x+3x)
4x=48 (subtract 10 from each side)
x=12 (divide each side by 4)
Therefore the son is 12 years old. The dad is 46 years old. The difference in their ages is 34.